# arXiv:2603.00996v1[cs.HC] 01 Mar 2026

### v1

@@ 1. Background
For the child-computer interaction community, this framing opens important design questions. Technologies can exacerbate the awareness-agency mismatch through fear-based content, algorithmic amplification of distress, and passive consumption patterns (kim2025privacynorm; radesky2022moralpanic; lukoff2018mindless). But technologies can also support sustainable care—through games that model civic engagement, tools that connect children to communities of practice, or platforms that make pathways visible and achievable (kim2024positech; parnes2025digitalinnovations; landesman2023letkids). The question is not whether children should encounter difficult realities but how technology might mediate that encounter in ways that sustain rather than deplete their capacity to care.

### v2

# Sustainable Care: Designing Technologies That Support Children’s Long-Term Engagement with Social Issues

@@ 1. Background
For the child-computer interaction community, this framing opens important design questions. Technologies can exacerbate the awareness-agency mismatch through fear-based content, algorithmic amplification of distress, and passive consumption patterns (kim2025privacynorm; radesky2022moralpanic; lukoff2018mindless; dangol2025genai). But technologies can also support sustainable care—through games that model civic engagement, tools that connect children to communities of practice, or platforms that make pathways visible and achievable (kim2024positech; parnes2025digitalinnovations; landesman2023letkids; Dangol et al., 2025). The question is not whether children should encounter difficult realities but how technology might mediate that encounter in ways that sustain rather than deplete their capacity to care.